2 masses m and m/2 are joined together by means of a light inextensible string,passing over a frictionless pulley.when the bigger mass is released,the small one will ascend with an acceleration g/n,where g is acceleration due to gravity.find n

Grade:11

1 Answers

Archita
8 years ago
m1=m/2          m2=m         a=g/n
 
a=[(m2-m1)/(m2+m1)]g
g/n=[(m-m/2)/(m+m/2)]g
on solving,
 
1/n=1/3
n=3
 
if u have any doubt ask in ans box.
approve if helpful.
this is a shortcut, if the ans printed in ur book is not this pls let me know so that i can work out this problem in the orginal method